How many gallons of cream containing 25% butter fat and milk containing 3 1/2% butter fat must be mixed to obtain 50 gallons of cream containing 12 1/2% butter fat. And also how much cream and how much milk to the nearest hundredth of a gallon must b mixed together to obtain 50 gallons of cream that contains 12.5% butterfat?
Answer by mananth(12270)

percent ----------------quantity
Cream A 25 ----------------x gallons
Milk 3.5 ----------------50-x gallons
Total 12.50% ----------------50

25x+3.5 (50-x)= 12.5*50
25x+175 -3.5x=625
25x-3.5 x=625-175
21.5x=450
/21.5
x=20.93 gallons 25.00%
29.07 gallons 3.50%